A reaction 𝐴 + 𝐵 → 𝑃 has been studied with the half-time method to determine a total order of a reaction and the rate constant. The experimental measurements determined the following characteristic equation:

log6𝑡!/#7 = 0.290618 + 0.500 ∗ log(𝐴$)

(a) Find the total order, and the rate constant of the reaction.

(b) State if and when one has to be cautious using the half-time method to determine the order or the rate constant of a reaction.

(c) State if the method can be used to determine orders of reaction with respect to individual reactants. If yes, briefly describe what needs to be done to determine these orders.


1
Expert's answer
2021-05-25T07:51:11-0400

(a)

a plot of log(t1/2) vs log(Ao) will have, slope = 1 - n

we have, 1-n = 1.200

n = - 0.200

==>

we have intercept , log {(2^n-1 -1)/k*(n-1) } = 0.228958

{(2^n-1 -1)/k*(n-1) } = 10^0.228958 = 1.694

0.4706/k = 1.694

k = 0.2778


(b)   

Usually, this method applies with rate law involving only one concentration term (when the rate law has the form r = k[A]n).


c) When order of reaction is half this method cannot be applied.

for n=1 , t1/2 is independent of [A]o.